OpenStudy (anonymous):

How to Solve r^2-10r+24=0

OpenStudy (anonymous):

r^2 - 10r + 25 - 1 = 0 (r - 5)^2 - 1 = 0 (r - 5 -1 )(r - 5 + 1) = 0 (r - 6)(r - 4) = 0